Based on the following information, what is the variance? State of Economy Probability of State of Economy Rate of Return if State Occurs Recession 0.29 − 9.70% Normal 0.40 11.20% Boom 0.31 22.20%

A) 0.12536 B) 0.02357 C) 0.01572 D) 0.03143 E) 0.07840
Answer #1

Expected return=Respective return*Respective probability

=(0.29*-9.7)+(0.4*11.2)+(0.31*22.2)=8.549%

probability Return probability*(Return-Expected Return)^2
0.29 -9.7 0.29*(-9.7-8.549)^2=96.5775403
0.4 11.2 0.4*(11.2-8.549)^2=2.8111204
0.31 22.2 0.31*(22.2-8.549)^2=57.7684383
Total=157.157099%

Standard deviation=[Total probability*(Return-Expected Return)^2/Total probability]^(1/2)

=(157.157099)^(1/2)

=12.54%(Approx)

Variance=Standard deviation^2

=0.01572 (Approx).